Which one of the following is a better buy: a large pizza with a 14-inch diameter for $12.00 or a medium pizza with a 7-inch diameter for $5.00?
Found 2 solutions by stanbon, oberobic:
Answer by stanbon(75887)

Which one of the following is a better buy: a large pizza with a 14-inch diameter for $12.00 or a medium pizza with a 7-inch diameter for $5.00?
-----
14" area: (pi)14^2 sq.in
Price per sq. in: [pi*14^2]/1200 == 0.51 cents/sq in
----
7" area: (pi)7^2 sq.in
Price per sq in: (pi*7^2)/500 = 0.31 cents/sq in
-----
Medium is cheaper.
Cheers,
Stan H.

Answer by oberobic(2304)

Area = pi * r^2
.
A pizza with a 7-inch diameter has a radius of 3.5 inches.
Area = pi*3.5^2 = 38.48 sq. in.
.
A pizza with a 14-inch diameter has a radius of 7 inches.
Area = pi*7^2 = 153.93 sq. in.
.
The smaller pizza costs $5 = 500 cents, so it costs 500/38.48 = about 13 cents per sq. in.
.
It costs $12 = 1200 cents, so it costs 12/153.93 = about 8 cents per sq in.
.
The larger pizza is a better buy in terms of the cost per sq. in.
.
Done.